# GST-UNet: Spatiotemporal Causal Inference with Time-Varying Confounders

### v1

@@ Abstract
Estimating causal effects from spatiotemporal data is a key challenge in fields such as public health, social policy, and environmental science, where controlled experiments are often infeasible. However, existing causal inference methods relying on observational data face significant limitations: they depend on strong structural assumptions to address spatiotemporal challenges – such as interference, spatial confounding, and temporal carryover effects – or fail to account for time-varying confounders. These confounders, influenced by past treatments and outcomes, can themselves shape future treatments and outcomes, creating feedback loops that complicate traditional adjustment strategies. To address these challenges, we introduce the GST-UNet (G-computation Spatio-Temporal UNet), a novel end-to-end neural network framework designed to estimate treatment effects in complex spatial and temporal settings. The GST-UNet leverages regression-based iterative G-computation to explicitly adjust for time-varying confounders, providing valid estimates of potential outcomes and treatment effects. To the best of our knowledge, the GST-UNet is the first neural model to account for complex, non-linear dynamics and time-varying confounders in spatiotemporal interventions. We demonstrate the effectiveness of the GST-UNet through extensive simulation studies and showcase its practical utility with a real-world analysis of the impact of wildfire smoke on respiratory hospitalizations during the 2018 California Camp Fire. Our results highlight the potential of GST-UNet to advance spatiotemporal causal inference across a wide range of policy-driven and scientific applications.
Spatiotemporal causal inference, time-varying confounders, G-computation, neural networks, treatment effects, observational data
@@ 1 Introduction
Environmental hazards, public health interventions, and large-scale socio-economic policies often require understanding complex cause-and-effect relationships across space and time (Reid et al., 2016b; Papadogeorgou et al., 2019; Song et al., 2020). For instance, when evaluating air quality interventions, policymakers need to assess how industrial regulations affect both immediate local health outcomes and broader regional impacts that evolve over time. These applications demand robust methods for estimating causal effects from observational spatiotemporal data.
However, spatiotemporal causal inference poses unique challenges. First, outcomes at each location are typically influenced by both local and neighboring covariates and interventions, leading to spatial confounding and interference. Second, observations exhibit strong temporal dependencies, with the effects of interventions “carrying over” in time. Finally, time-varying confounders–covariates that both influence and are influenced by past treatments and outcomes–create feedback loops that violate standard assumptions of independence over time. Consider, for example, how air quality policies are often implemented: governments may impose stricter regulations in response to poor air quality and high hospitalization rates, which in turn affect future air quality and health outcomes. Moreover, when estimating effects, we aim to understand the impact of different intervention sequences at specific locations given the observed history—a more challenging task than estimating average effects across space or time.
Current approaches to spatiotemporal causal inference either rely on restrictive modeling assumptions that may not capture real-world complexity, or employ flexible neural networks that are limited to single time points, independent time series, or settings without time-varying confounding (see Section 2 for a comprehensive overview). Further complicating matters, spatiotemporal settings often provide only a single realization of the process over time, rather than multiple independent time series.
To bridge this methodological gap, we propose GST-UNet (G-computation Spatiotemporal UNet), a novel end-to-end neural network framework designed to estimate treatment effects in complex spatiotemporal settings. GST-UNet addresses the key challenges of interference, spatial confounding, temporal carryover, and time-varying confounding by combining two powerful approaches: a U-Net-based encoder-decoder network for capturing spatial dependencies, and a regression-based, iterative G-computation procedure (Bang and Robins, 2005; Robins and Hernan, 2008) for adjusting for time-varying covariates. This integration enables valid estimation of potential outcomes while flexibly modeling complex dynamics across the spatial grid.
Our contributions are threefold: (1) We introduce the GST-UNet model architecture and provide theoretical foundations and implementation details for its iterative G-computation scheme—to the best of our knowledge, this is the first neural model to account for complex spatial dynamics and time-varying confounders in spatiotemporal interventions; (2) We demonstrate GST-UNet’s effectiveness through extensive simulation studies designed to showcase key spatiotemporal complexities and time-varying confounders; and (3) We illustrate its practical utility through a real-world analysis of wildfire smoke impacts on respiratory hospitalizations during the 2018 Camp Fire in California.
@@ 2 Related Work
We summarize the most relevant prior work here, with a more detailed discussion in Appendix A.
Classical Spatiotemporal Causal Inference. Earlier spatiotemporal causal inference methods (e.g., spatial econometrics (Anselin, 2013), difference-in-differences (Keele and Titiunik, 2015), and synthetic controls (Ben-Michael et al., 2022)) rely on strong assumptions (e.g., parallel trends) and often fail to address interference or time-varying confounders. More recent classical approaches, on the other hand, typically estimate average effects at the regional level or rely on structural and modeling assumptions that may not hold in real-world spatiotemporal contexts (Wang, 2021; Christiansen et al., 2022; Papadogeorgou et al., 2022; Zhang and Ning, 2023; Zhou et al., 2024).
Machine Learning for Spatiotemporal Modeling. Machine learning models (e.g., convolutional and recurrent networks-based methods (Shi et al., 2015; Zhang et al., 2017), or graph-based approaches (Li et al., 2017; Wu et al., 2019)) capture spatiotemporal patterns for prediction but lack formal causal adjustments.
Time Series Causal Inference. Time-series causal inference often uses recurrent or transformer-based methods (Bica et al., 2020; Seedat et al., 2022; Melnychuk et al., 2022) but assumes independent time series, ignoring potential interference effects. Although iterative G-computation (Bang and Robins, 2005; Robins and Hernan, 2008) or marginal structural models (Robins et al., 2000) can handle time-varying confounders, most ML extensions (Lim, 2018; Li et al., 2021; Hess et al., 2024) exclude interference or cross-series confounding.
Neural-Based Spatiotemporal Causal Inference. In the context of neural spatiotemporal models, Tec et al. (2023) integrate spatial representations for causal inference, accounting for spatial confounding and leveraging temporal data to train a UNet model. However, they do not address feedback effects from lagged or time-varying confounders. Most similar to our work, Ali et al. (2024) present a climate-focused model that shares certain architectural similarities but emphasizes prediction rather than adjusting for time-varying confounders, leaving causal identification concerns largely unaddressed.
@@ 3 Problem Formulation
Figure 1: Observational data (left) versus interventional data (right) for a horizon $\tau=2$ across multiple locations $(s,s^{\prime})$. Green arrows indicate temporal carryover, blue arrows show spatial confounding, and red arrows depict interference; dashed arrows denote time-varying confounding, and dashed circles represent unobserved variables at inference time. Under the intervention (right), treatments are set independently of confounders, and the full history is not observed for the entire horizon.
@@
Spatiotemporal Data. We model the observed data as random variables on a discrete spatial domain represented by an $N_{X}\times N_{Y}$ lattice: $\mathcal{S}=\{(i,j)\mid i\in[N_{X}],j\in[N_{Y}]\},$ where $[N]=\{1,\dots,N\}$ denotes the index set. Time is indexed by $t\in[T]$. At each spatial location $s=(i,j)$ at time $t$, we observe a tuple $(\mathbf{X}_{s,t},A_{s,t},Y_{s,t})$, where $A_{s,t}\in\{0,1\}$ represents a binary treatment (or intervention), $Y_{s,t}\in\mathbb{R}$ is a continuous outcome of interest, and $\mathbf{X}_{s,t}\in\mathbb{R}^{d_{X}}$ is a vector of time-varying covariates (e.g. local weather conditions, pollution levels, or socioeconomic indicators). Additionally, each location $s$ is associated with static features $V_{s}\in\mathbb{R}^{d_{v}}$ (e.g. geographical characteristics and socioeconomic indicators). While we focus on binary interventions for clarity, the methods generalize to more complex treatments. Conceptually, the data forms a 3D spatiotemporal tensor of size $T\times N_{X}\times N_{Y}$, though in practice, observations may be incomplete. Missing data can be accommodated using masking techniques during downstream modeling.
@@
To streamline notation, we use boldface symbols for random variables defined over the entire spatial domain. For $U\in\{X,A,Y\}$, let $\mathbf{U}_{t}$ denote its value at time $t$, and let $\mathbf{U}_{t:t+\tau}=(\mathbf{U}_{t},\dots,\mathbf{U}_{t+\tau})$ denote its value over a time interval. For a specific location $s$, we write $U_{s,t:t+\tau}=(U_{s,t},\dots,U_{s,t+\tau})$. The history up to time $t$ is denoted by $\mathbf{H}_{1:t}=(\mathbf{X}_{1:t},\mathbf{A}_{1:t-1},\mathbf{Y}_{1:t},\mathbf{V})$ for the entire spatial domain and $H_{s,1:t}=(X_{s,1:t},A_{s,1:t-1},Y_{s,1:t})$ for a specific location $s$. Specific instantiations of these random variables are denoted using lowercase letters (e.g., $u\in\{x,a,y,h\}$).
@@
Quantities of Interest. Our primary goal is to estimate location-specific Conditional Average Potential Outcomes (CAPOs) for a sequence of future spatiotemporal interventions, conditioned on observed history. Our approach builds on Rubin’s potential outcomes framework (Rubin, 1978; Robins and Hernan, 2008; Robins et al., 2000), which we extend to accommodate spatiotemporal settings. More concretely, we consider a future time horizon of length $\tau\geq 1$ and a predetermined interventional sequence $\mathbf{a}_{t:t+\tau-1}$ applied across the spatial domain starting at time $t$. Our goal is to estimate the potential outcomes at time $t+\tau$, denoted as $\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]$. In particular, we aim to compute:
@@
which represents the CAPOs at time $t+\tau+1$ under the given treatment sequence. Given two different interventional sequences $\mathbf{a}_{t:t+\tau}$ and $\mathbf{a}^{\prime}_{t:t+\tau}$, a related secondary goal is to estimate the location specific Conditional Average Treatment Effect (CATE), given by:
@@
Assumptions. Identification of CAPOs from observational data relies on standard causal inference assumptions. Additionally, our setup relies on observing only a single spatiotemporal outcome path. This prevents direct estimation of the CAPOs in Eq. (1) – which relies on an expectation over multiple data samples – without additional assumptions. We therefore introduce the following assumptions:
Assumption 3.1 (Standard Causal Inference Assumptions).
We assume the following properties hold: (Consistency) $\mathbf{Y}_{t+\tau}=\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]$ whenever the observed sequence of treatments $\mathbf{A}_{t:t+\tau-1}$ satisfies $\mathbf{A}_{t:t+\tau-1}=\mathbf{a}_{t:t+\tau-1}$;  (Positivity) $P(A_{s,t}=a_{s,t}\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t})>0$ for any $a_{s,t}\in\{0,1\}$ and feasible realization of history $\mathbf{h}_{1:t}$;   (Sequential Unconfoundedness) $\mathbf{Y}_{t+1:T}[\mathbf{a}_{t+1:T}]\perp\mathbf{A}_{t}\mid\mathbf{H}_{1:t}$, $\forall\mathbf{a}_{t+1:T}\in\{0,1\}^{T-t}$, i.e. at each time step $t$, the treatment assignment is independent of future potential outcomes.
Assumption 3.2 (Representation-Based Time Invariance).
@@
$p(\mathbf{X}_{t+1},\mathbf{Y}_{t+1}\mid\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})=z)$
$\quad=p(\mathbf{X}_{t^{\prime}+1},\mathbf{Y}_{t^{\prime}+1}\mid\phi(\mathbf{H}_{1:t^{\prime}},\mathbf{A}_{t^{\prime}})=z).$
@@
3.1 is standard in longitudinal causal inference settings (e.g., (Robins et al., 2000; Robins and Hernan, 2008; Bica et al., 2020; Li et al., 2021; Melnychuk et al., 2022; Hess et al., 2024)). 3.2 is specific to the single-time series setting, where pooling information across time is essential to enable estimation. We note that the single time-series setting frequently arises in causal inference, where assumptions such as stationarity or strict time homogeneity enable consistent estimation (Bojinov and Shephard, 2019; Papadogeorgou et al., 2022; Zhou et al., 2024). In contrast, our representation-based time invariance is weaker: rather than requiring $\mathbf{X}_{t},\mathbf{Y}_{t}$ themselves to have a time-invariant distribution, we only assume that, once the history is summarized by $\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})$, the transition to $(\mathbf{X}_{t+1},\mathbf{Y}_{t+1})$ follows a single shared mechanism. This approach aligns with modern time-series causal inference that learn time-invariant latent embeddings to pool information across time steps (Lim, 2018; Li et al., 2021; Hess et al., 2024), thus leveraging more data for a single, stable representation rather than time-dependent parameters.
Identification and G-Computation. We splice our single time series into multiple “prefixes” of varying lengths such that for each $t\in\{1,\dots,T-\tau\}$, we define the following
$\mathbf{P}_{t}^{\tau}\;=\;\bigl{(}\mathbf{X}_{1:t+\tau},\mathbf{A}_{1:t+\tau},\mathbf{Y}_{1:t+\tau},\mathbf{V}\bigr{)}.$
Under time-invariance (3.2), conditioning on a suitable embedding of $\mathbf{P}_{t}^{\tau}$ renders the distribution of $\mathbf{Y}_{t+\tau}$ independent of $t$. We will thus write expectations over the prefixes given history embeddings as
$\mathbb{E}_{\mathbf{P}}\!\bigl{[}\mathbf{Y}_{t+\tau}\;\big{|}\;\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})\bigr{]},$
where $t$ here only identifies the location in the series (i.e. the segment’s ending point) rather than implying a distinct distribution. Pooling across $t$ then supplies $T-\tau$ conditionally independent segments from a single time series, enabling regression-based methods to estimate future outcomes from (embedded) histories.

### v2

# GST-UNet: A Neural Framework for Spatiotemporal Causal Inference with Time-Varying Confounding

@@ Abstract
Estimating causal effects from spatiotemporal observational data is essential in public health, environmental science, and policy evaluation, where randomized experiments are often infeasible. Existing approaches, however, either rely on strong structural assumptions or fail to handle key challenges such as interference, spatial confounding, temporal carryover, and time-varying confounding—where covariates are influenced by past treatments and, in turn, affect future ones. We introduce the GST-UNet (G-computation Spatio-Temporal UNet), a theoretically grounded neural framework that combines a U-Net-based spatiotemporal encoder with regression-based iterative G-computation to estimate location-specific potential outcomes under complex intervention sequences. GST-UNet explicitly adjusts for time-varying confounders and captures non-linear spatial and temporal dependencies, enabling valid causal inference from a single observed trajectory in data-scarce settings. We validate its effectiveness in synthetic experiments and in a real-world analysis of wildfire smoke exposure and respiratory hospitalizations during the 2018 California Camp Fire. Together, these results position GST-UNet as a principled and ready-to-use framework for spatiotemporal causal inference, advancing reliable estimation in policy-relevant and scientific domains.
@@ 1 Introduction
Environmental hazards, public health interventions, and socio-economic policies often require understanding complex cause-and-effect relationships across space and time [34, 30, 41]. For instance, evaluating the health impacts of air quality regulations requires assessing how interventions influence both immediate outcomes and downstream effects across regions. Such applications demand robust tools for estimating causal effects from observational spatiotemporal data.
However, causal inference in spatiotemporal settings poses unique challenges. Outcomes are influenced not only by local covariates and interventions but also by those of neighboring regions (spatial confounding and interference). Effects may persist and accumulate over time (temporal carryover), and covariates often evolve in response to past interventions while simultaneously affecting future ones (time-varying confounding). For example, air quality regulations are often implemented in reaction to recent pollution levels and hospitalizations, which themselves shape future exposures and health outcomes—creating feedback loops that violate standard independence assumptions. These complexities induce bias in naive estimators and are especially challenging in single-trajectory settings, where replication across units or time is infeasible.
Existing approaches offer limited solutions: classical methods rely on rigid structural assumptions or user-defined exposure mappings, while recent neural models emphasize predictive accuracy over causal identification. Many assume independent time series or model only spatial correlations, leaving a gap in methods that can jointly address interference, temporal dependencies, and evolving confounding within a principled causal framework (see Section˜2).
To bridge this gap, we introduce GST-UNet (G-computation Spatio-Temporal UNet), a theoretically grounded neural framework for estimating location-specific potential outcomes in spatiotemporal settings with time-varying confounding. GST-UNet builds on formal identification and consistency results derived under a representation-based time-invariance assumption, showing how causal effects can be recovered from a single observed trajectory. We then instantiate this theory in a practical neural architecture: a U-Net encoder with ConvLSTM and attention modules coupled to an iterative G-computation procedure that performs recursive causal adjustment over time. To ensure stable estimation over long horizons, we design a curriculum-based training strategy that gradually refines recursive pseudo-outcomes, enabling effective learning even in data-scarce regimes. Unlike existing approaches, GST-UNet requires no user-specified structural models and can be directly deployed in real-world spatiotemporal applications.
Our contributions are threefold: (1) We develop the first unified framework that couples theoretical identification and consistency guarantees with an end-to-end neural implementation for spatiotemporal causal inference; (2) We demonstrate through controlled simulations that GST-UNet robustly handles interference, temporal carryover, and time-varying confounding; and (3) We illustrate its practical value via a real-world analysis of wildfire smoke exposure and respiratory hospitalizations during the 2018 California Camp Fire.
In summary, GST-UNet provides a principled and ready-to-use framework for causal inference from spatiotemporal data, combining formal guarantees with a flexible neural implementation. By abstracting away model-specific assumptions, GST-UNet makes spatiotemporal causal estimation both accessible and reliable for applied scientific and policy domains.
@@ 2 Related Work
We summarize the most relevant prior work here, with a more detailed discussion in Appendix˜A.
Classical Spatiotemporal Causal Inference. Early approaches (e.g., spatial econometrics [2], difference-in-differences [20], synthetic controls [4]) rely on strong assumptions such as parallel trends and no interference. More recent methods incorporate time-varying confounding using inverse propensity weighting (IPW) and marginal structural models [31, 49], but cannot address interference unless via user-specified exposure mappings or hyper-local assumptions [44, 11, 48]. As noted by Zhou et al. [49], the literature remains sparse, particularly in settings with rich feedback dynamics.
Machine Learning for Spatiotemporal Modeling. Deep learning models for prediction–e.g., CNNs and RNNs [40, 47], graph-based methods [25, 46], and video transformers [6, 27]–capture complex spatial-temporal patterns but do not incorporate causal adjustments, and thus cannot estimate counterfactuals or adjust for time-varying confounders.
Time Series Causal Inference. Causal methods for longitudinal data include marginal structural models [36], iterative G-computation [35], and recent ML-based extensions using recurrent networks, transformers, or meta-learners [7, 39, 28, 24, 16, 18]. However, these assume access to independent time series (e.g. across patients) and cannot model cross-unit interactions in spatiotemporal settings.
Neural-Based Spatiotemporal Causal Inference. Tec et al. [42] propose a UNet-based model that adjusts for non-local spatial confounding but focuses on static exposures and does not address interference or time-varying effects. Most similar to our work, [1] presents a climate-focused model that shares certain architectural similarities but emphasizes prediction rather than causal adjustment, leaving causal identification under time-varying confounding largely unaddressed.
Positioning of Our Work. Our work bridges these threads by uniting a theoretically grounded G-computation framework with a neural architecture for spatiotemporal data. Unlike prior time-series methods that assume independent units or spatial models that overlook confounding feedback, GST-UNet is the first end-to-end approach that (i) establishes identification and consistency under explicit assumptions for a single spatiotemporal trajectory, and (ii) implements this theory in a practical neural model capable of handling interference, spatial confounding, and time-varying dynamics.
@@ 3 Problem Formulation
Figure 1: Observational data (left) versus interventional data (right) for a horizon $\tau=2$ across multiple locations $(s,s^{\prime})$. Under the intervention (right), treatments are set independently of confounders, and the full history is not observed for the entire horizon.
@@
Spatiotemporal Data. We model observed data as random variables on a discrete spatial domain represented by an $N_{X}\times N_{Y}$ lattice: $\mathcal{S}=\{(i,j)\mid i\in[N_{X}],j\in[N_{Y}]\},$ where $N]=\{1,\dots,N\}$ denotes the index set. Time is indexed by $t\in[T]$. At each spatial location $s=(i,j)$ at time $t$, we observe a tuple $(\mathbf{X}_{s,t},A_{s,t},Y_{s,t})$, where $A_{s,t}\in\{0,1\}$ represents a binary treatment (or intervention), $Y_{s,t}\in\mathbb{R}$ is a continuous outcome of interest, and $\mathbf{X}_{s,t}\in\mathbb{R}^{d_{X}}$ is a vector of time-varying covariates (e.g. local weather conditions, pollution levels, or socioeconomic indicators). Additionally, each location $s$ is associated with static features $V_{s}\in\mathbb{R}^{d_{v}}$ (e.g. geographical characteristics and socioeconomic indicators). While we focus on binary interventions for clarity, the methods generalize to more complex treatments. Conceptually, the each variable forms a 3D spatiotemporal tensor of size $T\times N_{X}\times N_{Y}$, though in practice, observations may be incomplete. Missing data can be accommodated using masking techniques during downstream modeling.
@@
To streamline notation, we use boldface symbols for random variables defined over the entire spatial domain. For $U\in\{X,A,Y\}$, let $\mathbf{U}_{t}$ denote its value at time $t$, and let $\mathbf{U}_{t:t+\tau}=(\mathbf{U}_{t},\dots,\mathbf{U}_{t+\tau})$ denote its value over a time interval. For a specific location $s$, we write $U_{s,t:t+\tau}=(U_{s,t},\dots,U_{s,t+\tau})$. The history up to time $t$ is denoted by $\mathbf{H}_{1:t}=(\mathbf{X}_{1:t},\mathbf{A}_{1:t-1},\mathbf{Y}_{1:t},\mathbf{V})$ for the entire spatial domain and $H_{s,1:t}=(X_{s,1:t},A_{s,1:t-1},Y_{s,1:t},V_{s})$ for a specific location $s$. Specific instantiations of these random variables are denoted using lowercase letters (e.g., $u\in\{x,a,y,h\}$).
@@
Quantities of Interest. Our primary goal is to estimate location-specific Conditional Average Potential Outcomes (CAPOs) for a sequence of future spatiotemporal interventions, conditioned on observed history. Our approach builds on Rubin’s potential outcomes framework [38, 35, 36], which we extend to accommodate spatiotemporal settings. More concretely, we consider a future time horizon of length $\tau\geq 1$ and a predetermined interventional sequence $\mathbf{a}_{t:t+\tau-1}$ applied across the spatial domain starting at time $t$. Our goal is to estimate the potential outcomes at time $t+\tau$, denoted as $\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]$. In particular, we aim to compute:
@@
which represents the CAPOs at time $t+\tau$ under the given treatment sequence. Given two different interventional sequences $\mathbf{a}_{t:t+\tau}$ and $\mathbf{a}^{\prime}_{t:t+\tau}$, a related secondary goal is to estimate the location specific Conditional Average Treatment Effect (CATE), given by:
@@
Prefix Data in a Single Spatiotemporal Chain. The conditional expectations defining the CAPOs in Eq. (1) cannot be directly estimated from a single observed spatiotemporal realization, since the empirical averages would contain only one sample of each future outcome $\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]$. To obtain a workable regression-based estimator, we therefore reorganize the single observed trajectory into overlapping "prefixes" of varying lengths. For each $t\in\{1,\dots,T-\tau\}$, we define
$\mathbf{P}_{t}^{\tau}\;=\;\bigl(\mathbf{X}_{1:t+\tau},\mathbf{A}_{1:t+\tau},\mathbf{Y}_{1:t+\tau},\mathbf{V}\bigr),$
which represents the observed history up to time $t+\tau$ along with all covariates, treatments, and outcomes. When $T\!\gg\!\tau$, this construction yields $T-\tau$ segments that partially overlap in time, providing additional training samples in this intrinsically data-scarce, single-chain setting.
However, these prefixes are not independent: successive segments share overlapping histories, so standard i.i.d. assumptions do not apply. In the next section, we introduce conditions under which these prefixes can be treated as conditionally exchangeable given an appropriate learned embedding. This enables regression-based estimation of CAPOs by pooling information across time without violating the dependence structure of the original process.
4 Identification and Estimation of CAPOs in Spatiotemporal Settings
Identification of CAPOs from observational data relies on standard causal inference assumptions. In our setting, these must be complemented by additional structure to handle the fact that we observe only a single spatiotemporal trajectory. Building on the prefix construction introduced above, we impose conditions that render these overlapping segments conditionally exchangeable, enabling principled pooling of information across time.
Assumption 1 (Causal Inference Assumptions).
We assume: (Consistency) $\mathbf{Y}_{t+\tau}=\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]$ whenever the observed sequence of treatments $\mathbf{A}_{t:t+\tau-1}$ satisfies $\mathbf{A}_{t:t+\tau-1}=\mathbf{a}_{t:t+\tau-1}$;  (Positivity) $P(A_{s,t}=a_{s,t}\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t})>0$ for any $a_{s,t}\in\{0,1\}$ and feasible realization of history $\mathbf{h}_{1:t}$;  (Sequential Unconfoundedness) $\mathbf{Y}_{t+1:T}[\mathbf{a}_{t+1:T}]\perp\mathbf{A}_{t}\mid\mathbf{H}_{1:t}$, $\forall\mathbf{a}_{t+1:T}\in\{0,1\}^{T-t}$, i.e. at each time step $t$, the treatment assignment is independent of future potential outcomes.
Assumption 2 (Representation-Based Time Invariance).
@@ 4 Identification and Estimation of CAPOs in Spatiotemporal Settings
$p(\mathbf{X}_{t+1},\mathbf{Y}_{t+1}\mid\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})=z)=p(\mathbf{X}_{t^{\prime}+1},\mathbf{Y}_{t^{\prime}+1}\mid\phi(\mathbf{H}_{1:t^{\prime}},\mathbf{A}_{t^{\prime}})=z).$
@@
Assumption˜1 is a standard set of requirements in longitudinal causal inference settings (e.g., [36, 35, 7, 24, 28, 18]). Assumption˜2 is specific to the single-time series setting, where pooling information across time is essential to enable estimation. We note that the single time-series setting frequently arises in causal inference, where assumptions such as stationarity or strict time homogeneity enable consistent estimation [8, 31, 49]. In contrast, our representation-based time invariance is weaker: rather than requiring $\mathbf{X}_{t},\mathbf{Y}_{t}$ themselves to have a time-invariant distribution, we only assume that, once the history is summarized by $\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})$, the transition to $(\mathbf{X}_{t+1},\mathbf{Y}_{t+1})$ follows a single shared mechanism. This approach aligns with modern time-series causal inference that learn time-invariant latent embeddings to pool information across time steps [26, 24, 18], thus leveraging more data for a single, stable representation rather than time-dependent parameters.
Under Assumption˜2, conditioning on $\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})$ removes explicit dependence on $t$, such that
$\mathbb{E}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}\mid\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})]$
represents a shared conditional expectation across all prefix segments. In this view, $t$ indexes the segment’s position rather than a distinct distribution. Pooling over $t$ thus yields $T-\tau$ approximately exchangeable segments from a single trajectory, enabling regression-based estimation of future outcomes from embedded histories.
4.1 Identification via Representation-Based G-Computation